OpenStudy (anonymous):

The slope of a line is 1, and the y-intercept is -1. What is the equation of the line written in slope-intercept form? y = x - 1 y = 1 - x y = -x - 1

OpenStudy (anonymous):

Slope intercept form is: y=mx+b where mx is slope and b is y intercept.

OpenStudy (gabylovesyou):

Slope-Intercept Form: y= mx +b y = 1x + -1 y = x - 1 :)

OpenStudy (anonymous):

Use your slope intercept form and plug in your values. M is for your slope, which in this case is 1 and b will be your y value, which is -1. When you plug everything in its y=x-1